Calculate the force of gravity of the husband and wife, the husband has a mass of 85kg and the wife has a mass of 60kg. They are 0.75m apart.


1
Expert's answer
2021-12-07T09:55:10-0500
"F=\dfrac{Gm_hm_w}{d^2},""F=\dfrac{6.67\times10^{-11}\ \dfrac{N\times m^2}{kg^2}\times85\ kg\times60\ kg}{(0.75\ m)^2}=6.05\times10^{-7}\ N."